OpenStudy (heyitslizzy13):

can anyone teach me how to simplify numbers with fractions as exponents??? please! :)

OpenStudy (anonymous):

you got an example?

OpenStudy (heyitslizzy13):

i just want to know the basics so I can apply it to the question (81x^2/y)^3/4

OpenStudy (anonymous):

\[\huge b^{\frac{n}{m}}=\sqrt[m]{b^n}\]

OpenStudy (anonymous):

\[81^{\frac{3}{4}}=\sqrt[4]{81^3}=3^3=27\]

OpenStudy (heyitslizzy13):

so is that the same for every question?

OpenStudy (anonymous):

same idea if you have variables you just multipiy the exponents, so \[\huge (x^2)^{\frac{3}{4}}=x^{\frac{3}{2}}\]

OpenStudy (heyitslizzy13):

okay! thanks so much!! (: